# Almost Sure Convergence for the Last Iterate of Stochastic Gradient Descent Schemes

### v2

@@ Abstract.
We study the almost sure convergence rate for the last iterate of stochastic gradient descent (SGD) and stochastic heavy ball (SHB) in the parametric setting when the objective function $F$ is globally convex or non-convex whose gradient is $\gamma$-Hölder. Using only discrete Gronwall’s inequality without Robbins-Siegmund theorem, we recover results for both SGD and SHB: $\min_{s\leq t}\|\nabla F(w_{s})\|^{2}=o(t^{p-1})$ for non-convex objectives and $F(w_{t})-F_{*}=o(t^{2\gamma/(1+\gamma)\cdot\max(p-1,-2p+1)-\epsilon})$ for $\beta\in(0,1)$ and $\min_{s\leq t}F(w_{s})-F_{*}=o(t^{p-1})$ almost surely for convex objectives. In addition, we proved that SHB with constant momentum parameter $\beta\in(0,1)$ attains a convergence rate of $F(w_{t})-F_{*}=O(t^{\max(p-1,-2p+1)}\log^{2}\frac{t}{\delta})$ with probability at least $1-\delta$ when $F$ is convex and $\gamma=1$ and step size $\alpha_{t}=\Theta(t^{-p})$ with $p\in(\frac{1}{2},1)$.
@@ 1. Introduction
Another recent work regarding almost sure convergence rate for SGD without employing the Robbins-Siegmund theorem, is done by Nguyen et al [11]. The authors used a weaker, super-martingale convergence theorem [11, lemma 5] to prove $w_{t}\rightarrow w_{*}$ a.s.-$\mathbb{P}$ and $\mathbb{E}\|w_{t}-w_{*}\|^{2}=o(t^{-1})$ in the strong convex case [11, thorem 1,2]. Such a rate is indeed optimal according to [1] since if $F$ is convex with $\nabla F$ Lipschitz, by lemma B.4, we get $F(w_{t})-F_{*}\lesssim\|w_{t}-w_{*}\|^{2}=o(t^{-1})$.
@@ 1.2. Acknowledgment
I would like to thank my advisor, Sunder Sethuraman, for discussions and support throughout the writing process of this manuscript.
@@ 2.1. Notation and Assumptions
The assumption that $\nabla\ell$ is only $\gamma$-Hölder is used in [8, assumption 1], whose application is also mentioned in [8]. This is a weakening of the $\nabla\ell$ being Lipschitz used in [9, assumption 1]. However, thanks to a slight modification in the proof of Garrigos and Gower [4, lemma 2.25] (see proposition B.2), any function $f$ that is $(\gamma,L)$-smooth still enjoys the property
@@
such an assumption implies convergence almost surely of the empirical average on the right hand side to $\mathbb{E}_{\rho}[\nabla\ell(Z,w_{t})]=\nabla\mathbb{E}_{\rho}[\ell(Z,w_{t})]=\nabla F(w_{t})$ as the number of data points $n\uparrow\infty$. This suggests that (2.2) approximates the deterministic GD [4, algorithm 3.2]. In addition, there are important consequences. The first consequence is due to proposition B.1: $F$ inherits the properties in assumption 2.1 from the function $\ell(z,\cdot)$. Secondly, if the joint distribution of $(Z_{t},w_{t})$ is the product measure $\rho\,dz\otimes\mu_{t}$, then
@@
The above assumption is called the ABC condition, originally proposed by Khaled and Richtárik in [6], and is used by Liu and Yuan in [9, assumption 4] when $F$ is $(1,L)$-smooth, i.e. $\gamma=1$. It is said to be “the weakest assumption” for analysis of SGD in the non-convex setting [9, remark 1]. By lemma B.5, the first bounding term is a consequence of $(\gamma,L)$-smoothness and convexity:
@@
Assumption 2.3 is also used in [8]. Such an assumption can be satisfied when $\pazocal{Z}$ is bounded. See discussions in [8]. Note that by lemma B.6.3b, if $\ell(z,\cdot)$ is convex and assumption 2.1 holds, then the ‘noise’ $\|\delta m_{t}\|^{2}$ defined in (1.2) is bounded by a scalar multiple of $\|w_{t}-w_{*}\|^{2\gamma}+1$. This assumption will only be used to produce probability estimates for the convergence rate of SHB iterates.
@@ 2.2. Main Result 1: Almost Sure Convergence Rate
Let $Y_{t},Z_{t}$ be a sequence of non-negative random variables adapted to a filtration $\mathscr{F}_{t}$ and $c_{1},c_{2}\in\mathbb{R}_{+}$ be constants. Let $\alpha_{t}=\Theta(t^{-p})$ where $p\in(\frac{1}{1+\gamma},1)$ and $\gamma\in(0,1]$. If both of the following holds
@@
Compared to [9, lemma 13] and [13, theorem 1], we require a stronger condition, namely condition 2 where the expectation of the sum is required to be finite.
@@
      $F(w_{t})-F_{*}=o(t^{r_{\gamma}\max(p-1,1-(1+\gamma)p)+\epsilon})$ a.s.-$\mathbb{P}$ where $r_{\gamma}=\begin{cases}\frac{2\gamma}{1+\gamma}&\beta\in(0,1)\\ 1&\beta=0.\end{cases}$.
@@
The slowdown factor $r_{\gamma}$ due to the smoothness parameter $\gamma$ appears only with momentum $\beta>0$ and $\gamma<1$. When $\gamma<1$, intuitively, the algorithm is slowed down as can be seen in the last term of (2.1). The fact that the slowdown factor $r_{\gamma}$ is uniform for all $\beta>0$ is interesting as empirical evidence showed that larger $\beta$ contributes to more slowdown [3]. There are also other works on non-acceleration of heavy ball [5]. We do not quite understand at the moment of these phenomena and how it ties to the slowdown factor $r_{\gamma}$. In addition, it is counter-intuitive since momentum is supposed to make convergence faster when the gradient is small. Examining (3.13), when there is momentum, the gradient $\|\nabla F(w_{t}+\frac{\beta}{1-\beta}(w_{t}-w_{t-1}))\|^{2}$ appears as an additive term in the upper estimate for $F(w_{t})-F_{*}$ and lemma B.4 suggests that for $\gamma<1$, this slows down the rate.
@@ 2.3. Main Result 2: Convergence Rate with High Probability
From the previous paragraph, applying martingale concentration inequalities on $\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ requires controlling the noise $\delta m_{t}$. By definition, $\delta m_{t}$ is comprised of $\nabla F(w_{t})$ and $\nabla\ell(Z_{t},w_{t})$, which by lemma B.4, is controlled by $\|w_{t}-w_{*}\|$. Following [8], in proposition 4.5, we show that up to time $T>0$, $\|w_{t}-w_{*}\|$ is at most $\log(T/\delta)$ with probability $1-\delta$. Since $\alpha_{t}$ decays with order $O(t^{-p})$, the martingale difference $\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ is of the order $\alpha_{t}$ up to a logarithmic factor.
@@ 3.1. Stochastic Gradient Descent
Due to $(\gamma,L)$-smoothness of $F$ and assumption 2.2, by proposition B.2,
@@
Since $\alpha_{t}\in\ell^{2}(\mathbb{N})$, by proposition 3.1, the third term in the above estimate is summable. We may apply proposition 2.4 and conclude $\sum\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]<\infty$
@@
From proposition 3.1, we may conclude that $\mathbb{P}(\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty)=1$ and if $F$ is convex, $\mathbb{P}[\sum\alpha_{t}\,(F(w_{t})-F_{*})]=1$. Intuitively, this suggests that if $\sum\alpha_{t}=\infty$, then $\|\nabla F(w_{t})\|^{2}\rightarrow 0$ almost surely (resp. for the convex case $F(w_{t})-F_{*}\rightarrow 0$ almost surely). Thanks to [12, theorem 2], such intuition holds in our case. Note that [12, theorem 2] holds without using any martingale convergence theorem and that the bounded variance assumption can be replaced with assumption 2.2 as shown below.
@@
      If $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then $F(w_{t})-F_{*}=o(t^{\max(p-1,1-(1+\gamma)p)+\epsilon})$ a.s.-$\mathbb{P}$.
@@
Since proposition 3.1 asserts $\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty$ almost surely, by lemma B.3, it suffices to show that the second term on the right hand side has finite $L^{2}$ norm. Indeed, if this is the case, then $\sum_{i=s}^{t}\alpha_{i}\delta m_{i}<\infty$ almost surely. By proposition B.7 and assumption 2.2,
@@
For the convex case, we know that $F$ has no saddle points and all minima are global. Therefore, the set of critical points $\pazocal{C}=\{w\in\pazocal{W}:\|\nabla F(w)\|^{2}=0\}$ is the set of global minima. Therefore, $\|\nabla F(w_{t})\|^{2}\rightarrow 0$ implies $w_{t}\rightarrow w$ where $w\in\pazocal{C}$. By continuity of $F$, we get $F(w_{t})\rightarrow F(w)=F_{*}$. Moreover, defining $Y_{t}=F(w_{t})-F_{*}$, applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{s}]$ on (3.2), dropping the $\alpha_{t}\,\|\nabla F(w_{t})\|^{2}$ term, and applying lemma B.4 yields an inequality of the form
@@ 3.2. Stochastic Heavy Ball
      If $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then $F(w_{t})-F_{*}=o(t^{r_{\gamma}\max(p-1,1-(1+\gamma)p)+\epsilon})$ a.s.-$\mathbb{P}$ where
      $r_{\gamma}=\begin{cases}\frac{2\gamma}{1+\gamma}&\beta\in(0,1)\\ 1&\beta=0.\end{cases}$
@@
The same argument as in corollary 3.3 can be used. Both claims follow from (3.7) and (3.8) along with the first and second conclusion in proposition 3.4. Let $Y_{t}=F(z_{t})-F_{*}+\|v_{t}\|^{2}$. Applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{s}]$ on (3.11) and dropping the $\alpha_{t}\|\nabla F(z_{t})\|^{2}$ term yields
@@
Proposition 3.4 shows that all conditions in lemma 2.5 holds. Estimating $F(w_{t})-F_{*}$ using $(\gamma,L)$-smoothness and applying lemma B.4 yields
@@ 4. Convergence Rate with High Probability for Convex Objective
We will assume throughout this section that $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, which by proposition B.1, makes $F$ also convex. It will be convenient to define once and for all the following constants and notations. Let $a_{1}$ and $a_{2}$ be the constants defined in lemma B.6. Also, define the following
@@ 4.1. Distance from a Minimizer with High Probability
Consider (3.6) for the iterates of $v_{t}$. Take the squared-norm. Applying lemma B.6.3a yields
@@
Using the notation in (4.1), convexity and lemma B.6.3a yields
@@
Also, we have $\mathbb{E}[\xi_{t}\,|\,\mathscr{F}_{t}]=0$. We now show that we may bound $\sum_{k=1}^{t}\xi_{k}-\mathbb{E}[\xi_{k}\,|\,\mathscr{F}_{k}]$ using [8, lemma 21]. By lemma B.6.3b, lemma 4.1, and (4.9),
@@
As for the conditional variance, we apply Cauchy-Schwarz, lemma B.6.3d, and lemma 4.1 to get

### v3

# Convergence Rate for the Last Iterate of Stochastic Gradient Descent Schemes

@@ Abstract.
We study the convergence rate for the last iterate of stochastic gradient descent (SGD) and stochastic heavy ball (SHB) in the parametric setting when the objective function $F$ is globally convex or non-convex whose gradient is $\gamma$-Hölder. Using only discrete Gronwall’s inequality without Robbins-Siegmund theorem, we recover results for both SGD and SHB: $\min_{s\leq t}\|\nabla F(w_{s})\|^{2}=o(t^{p-1})$ for non-convex objectives and $F(w_{\tau\wedge t})-F_{*}=o(t^{2\gamma/(1+\gamma)\cdot\max(p-1,-2p+1)-\epsilon})$ for $\beta\in(0,1)$, $\tau:=\inf\{t>0:F(w_{t})=F_{*}\}$, and $\min_{s\leq t}F(w_{s})-F_{*}=o(t^{p-1})$ for convex objectives $F$ whose minimum is $F_{*}$. In addition, we proved that SHB with constant momentum parameter $\beta\in(0,1)$ attains a convergence rate of $F(w_{t})-F_{*}=O(t^{\max(p-1,-2p+1)}\log^{2}\frac{t}{\delta})$ with probability at least $1-\delta$ when $F$ is convex and $\gamma=1$ and step size $\alpha_{t}=\Theta(t^{-p})$ with $p\in(\frac{1}{2},1)$.
@@ 1. Introduction
Another recent work regarding almost sure convergence rate for SGD without employing the Robbins-Siegmund theorem, is done by Nguyen et al [11]. The authors used a weaker, super-martingale convergence theorem [11, lemma 5] to prove $w_{t}\rightarrow w_{*}$ a.s.-$\mathbb{P}$ and $\mathbb{E}\|w_{t}-w_{*}\|^{2}=o(t^{-1})$ in the strong convex case [11, thorem 1,2]. Such a rate is indeed optimal according to [1] since if $F$ is convex with $\nabla F$ Lipschitz, by lemma C.4, we get $F(w_{t})-F_{*}\lesssim\|w_{t}-w_{*}\|^{2}=o(t^{-1})$.
@@ 1.2. Acknowledgment
I would like to thank Prof. Sunder Sethuraman, for discussions and support throughout the writing process of this manuscript.
@@ 2.1. Notation and Assumptions
The assumption that $\nabla\ell$ is only $\gamma$-Hölder is used in [8, assumption 1], whose application is also mentioned in [8]. This is a weakening of the $\nabla\ell$ being Lipschitz used in [9, assumption 1]. However, thanks to a slight modification in the proof of Garrigos and Gower [4, lemma 2.25] (see proposition C.2), any function $f$ that is $(\gamma,L)$-smooth still enjoys the property
@@
such an assumption implies convergence almost surely of the empirical average on the right hand side to $\mathbb{E}_{\rho}[\nabla\ell(Z,w_{t})]=\nabla\mathbb{E}_{\rho}[\ell(Z,w_{t})]=\nabla F(w_{t})$ as the number of data points $n\uparrow\infty$. This suggests that (2.2) approximates the deterministic GD [4, algorithm 3.2]. In addition, there are important consequences. The first consequence is due to proposition C.1: $F$ inherits the properties in assumption 2.1 from the function $\ell(z,\cdot)$. Secondly, if the joint distribution of $(Z_{t},w_{t})$ is the product measure $\rho\,dz\otimes\mu_{t}$, then
@@
The above assumption is called the ABC condition, originally proposed by Khaled and Richtárik in [6], and is used by Liu and Yuan in [9, assumption 4] when $F$ is $(1,L)$-smooth, i.e. $\gamma=1$. It is said to be “the weakest assumption” for analysis of SGD in the non-convex setting [9, remark 1]. By lemma C.5, the first bounding term is a consequence of $(\gamma,L)$-smoothness and convexity:
@@
Assumption 2.3 is also used in [8]. Such an assumption can be satisfied when $\pazocal{Z}$ is bounded. See discussions in [8]. Note that by lemma C.6.3b, if $\ell(z,\cdot)$ is convex and assumption 2.1 holds, then the ‘noise’ $\|\delta m_{t}\|^{2}$ defined in (1.2) is bounded by a scalar multiple of $\|w_{t}-w_{*}\|^{2\gamma}+1$. This assumption will only be used to produce probability estimates for the convergence rate of SHB iterates.
@@ 2.2. Main Result 1: Almost Sure Convergence Rate
Let $Y_{t},Z_{t}$ be a sequence of positive and non-negative random variables respectively, adapted to a filtration $\mathscr{F}_{t}$ and $c_{1},c_{2}\in\mathbb{R}_{+}$ be constants. Let $\alpha_{t}=\Theta(t^{-p})$ where $p\in(\frac{1}{1+\gamma},1)$ and $\gamma\in(0,1]$. If both of the following holds
@@
Compared to [9, lemma 19] and [13, theorem 1], we require $Y_{t}>0$ and item 2 where the expectation of the sum is required to be finite. The application of lemma 2.5 will be for convex cost $F$ and $Y_{t}=F(w_{t})-F_{*}$. Due to the requirement $Y_{t}>0$, instead of stating general convergence rate for $F(w_{t})-F_{*}$, we only make statements about the convergence rate of $F(w_{\tau\wedge t})-F_{*}$ where $\tau:=\inf\{t>0:F(w_{t})=F_{*}\}$.
@@
      Let $\tau:=\inf\{t>0:F(w_{t})=F_{*}\}$. We have
      $F(w_{\tau\wedge t})-F_{*}=o(t^{r_{\gamma}\max(p-1,1-(1+\gamma)p)+\epsilon})~~\text{where}~~r_{\gamma}=\begin{cases}\frac{2\gamma}{1+\gamma}&\beta\in(0,1)\\ 1&\beta=0\end{cases}.$
@@
The slowdown factor $r_{\gamma}$ due to the smoothness parameter $\gamma$ appears only with momentum $\beta>0$ and $\gamma<1$. When $\gamma<1$, intuitively, the algorithm is slowed down as can be seen in the last term of (2.1). The fact that the slowdown factor $r_{\gamma}$ is uniform for all $\beta>0$ is interesting as empirical evidence showed that larger $\beta$ contributes to more slowdown [3]. There are also other works on non-acceleration of heavy ball [5]. We do not quite understand at the moment of these phenomena and how it ties to the slowdown factor $r_{\gamma}$. In addition, it is counter-intuitive since momentum is supposed to make convergence faster when the gradient is small. Examining (3.13), when there is momentum, the gradient $\|\nabla F(w_{t}+\frac{\beta}{1-\beta}(w_{t}-w_{t-1}))\|^{2}$ appears as an additive term in the upper estimate for $F(w_{t})-F_{*}$ and lemma C.4 suggests that for $\gamma<1$, this slows down the rate.
@@ 2.3. Main Result 2: Convergence Rate with High Probability
From the previous paragraph, applying martingale concentration inequalities on $\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ requires controlling the noise $\delta m_{t}$. By definition, $\delta m_{t}$ is comprised of $\nabla F(w_{t})$ and $\nabla\ell(Z_{t},w_{t})$, which by lemma C.4, is controlled by $\|w_{t}-w_{*}\|$. Following [8], in proposition 4.5, we show that up to time $T>0$, $\|w_{t}-w_{*}\|$ is at most $\log(T/\delta)$ with probability $1-\delta$. Since $\alpha_{t}$ decays with order $O(t^{-p})$, the martingale difference $\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ is of the order $\alpha_{t}$ up to a logarithmic factor.
@@ 3.1. Stochastic Gradient Descent
Due to $(\gamma,L)$-smoothness of $F$ and assumption 2.2, by proposition C.2,
@@
Since $\alpha_{t}\in\ell^{2}(\mathbb{N})$, by proposition 3.1, the third term in the above estimate is summable. We may apply proposition 2.4 and conclude $\sum\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]<\infty$. ∎
@@
From proposition 3.1, we may conclude that $\mathbb{P}(\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty)=1$ and if $F$ is convex, $\mathbb{P}[\sum\alpha_{t}\,(F(w_{t})-F_{*})]=1$. Intuitively, this suggests that if $\sum\alpha_{t}=\infty$, then $\|\nabla F(w_{t})\|^{2}\rightarrow 0$ almost surely (resp. for the convex case $F(w_{t})-F_{*}\rightarrow 0$ almost surely). Thanks to [12, theorem 2], such intuition holds in our case.
@@
      Let $\tau$ be defined as in theorem 2.7. If $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then almost surely,
      $F(w_{\tau\wedge t})-F_{*}=o(t^{\max(p-1,1-(1+\gamma)p)+\epsilon}).$
@@
Since proposition 3.1 asserts $\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty$ almost surely, by lemma C.3, it suffices to show that the second term on the right hand side has finite $L^{2}$ norm. Indeed, if this is the case, then $\sum_{i=s}^{t}\alpha_{i}\delta m_{i}<\infty$ almost surely. By proposition C.7 and assumption 2.2,
@@
Part 2.a. For the convex case, we know that $F$ has no saddle points and all minima are global. Therefore, the set of critical points $\pazocal{C}=\{w\in\pazocal{W}:\|\nabla F(w)\|^{2}=0\}$ is the set of global minima. Therefore, $\|\nabla F(w_{t})\|^{2}\rightarrow 0$ implies $w_{t}\rightarrow w$ where $w\in\pazocal{C}$. By continuity of $F$, we get $F(w_{t})\rightarrow F(w)=F_{*}$.
Part 2.b. For a realization $\omega\in\Omega$, if $\tau(\omega)<\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F_{*}$ for $t\geq\tau(\omega)$. The claim then holds trivially. If $\tau(\omega)=\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F(w_{t}(\omega))$ for all $t>0$. Defining $Y_{t}=F(w_{t})-F_{*}$, applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{s}]$ on (3.2), dropping the $\alpha_{t}\,\|\nabla F(w_{t})\|^{2}$ term, and applying lemma C.4 yields an inequality of the form
@@ 3.2. Stochastic Heavy Ball
      Let $\tau$ be defined as in theorem 2.7. If $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then almost surely
      $F(w_{\tau\wedge t})-F_{*}=o(t^{r_{\gamma}\max(p-1,1-(1+\gamma)p)+\epsilon})~~\text{where}~~r_{\gamma}=\begin{cases}\frac{2\gamma}{1+\gamma}&\beta\in(0,1)\\ 1&\beta=0\end{cases}.$
@@
The same argument as in corollary 3.3 can be used. For a realization $\omega\in\Omega$, if $\tau(\omega)<\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F_{*}$ for all $t\geq\tau(\omega)$ and the claim holds trivially. Otherwise, if $\tau(\omega)=\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F(w_{t}(\omega))$ so that the argument in the next paragraph can be used.
Both claims follow from (3.7) and (3.8) along with the first and second conclusion in proposition 3.4. Let $Y_{t}=F(z_{t})-F_{*}+\|v_{t}\|^{2}$. Applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{s}]$ on (3.11) and dropping the $\alpha_{t}\|\nabla F(z_{t})\|^{2}$ term yields
@@
Proposition 3.4 shows that all conditions in lemma 2.5 holds. Estimating $F(w_{t})-F_{*}$ using $(\gamma,L)$-smoothness and applying lemma C.4 yields
@@ 4. Convergence Rate with High Probability for Convex Objective
We will assume throughout this section that $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, which by proposition C.1, makes $F$ also convex. It will be convenient to define once and for all the following constants and notations. Let $a_{1}$ and $a_{2}$ be the constants defined in lemma C.6. Also, define the following
@@ 4.1. Distance from a Minimizer with High Probability
Consider (3.6) for the iterates of $v_{t}$. Take the squared-norm. Applying lemma C.6.3a yields
@@
Using the notation in (4.1), convexity and lemma C.6.3a yields
@@
Also, we have $\mathbb{E}[\xi_{t}\,|\,\mathscr{F}_{t}]=0$. We now show that we may bound $\sum_{k=1}^{t}\xi_{k}-\mathbb{E}[\xi_{k}\,|\,\mathscr{F}_{k}]$ using [8, lemma 21]. By lemma C.6.3b, lemma 4.1, and (4.9),
@@
As for the conditional variance, we apply Cauchy-Schwarz, lemma C.6.3d, and lemma 4.1 to get